# Spectraformer: A Unified Random Feature Framework for Transformer

### v3

@@ 2.2 Linearizing Kernelized Attention via Random Features
$\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$ (7)
@@
All that is left is to retrieve the feature map $\phi$, the process of which is called kernel approximation (see Section 2.3), which corresponds to the kernels we specify. There are many popular kernels from which we can choose, but typically, we want to approximate the softmax kernel in Equation 6. However, since the softmax relates to the $\mathcal{K}_{RBF}(\bm{x},\bm{y})=\exp(-\frac{||\bm{x}-\bm{y}||^{2}}{2})$ via the following equation:
$\mathcal{K}_{softmax}(\bm{x},\bm{y})=\exp(\frac{||\bm{x}||^{2}}{2})\mathcal{K}_{RBF}(\bm{x},\bm{y})\exp(\frac{||\bm{y}||^{2}}{2})$ (8)
@@
Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown in Tsai et al. (2019). Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4. Notwithstanding, we first introduce the main kernel approximation technique: random features.
2.3 Random Features for Kernel Approximation
Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi & Recht, 2007). Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem (Rudin, 1990) stated as follows:
@@ 2.3 Random Features for Kernel Approximation
inline]define unambiguously the dimension of x and phi here, also the dimension before and after approximation of the features map
$\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$ (9)
@@
$\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big{[}\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{\!\top}_{s}\bm{x})]^{\!\top}\end{array}$ (10)
2.4 Random Features for Kernel Learning
Previous random feature techniques are robust in approximating kernels for specific learning problems. However, kernels may be chosen using heuristics and by convention from a popular subset. The kernel is very much a hyperparameter. This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf & Smola, 2018). Tsai et al. (2019) has shown experimental results that this is also true in the context of attention in Transformer. Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique. Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson & Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022). Kernel learning is achieved by making the weight matrix $\bm{W}$ learnable where $\bm{\omega}_{i}$ is the output of a function parameterizing $p(\bm{\omega})$.
The kernelized attention via random features discussed here is a popular approach, however it is by no means the only one. A more detailed discussion is offered in related work in Section A.2. Specifically, Section A.2.1 provides alternative kernelized attention formulations and Section A.2.2 provides alternative kernel approximations to random features.
3 Spectraformer
@@
$\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big{[}a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$ (11)
Spectraformer is shown in Figure 1. The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\bm{v}$ and transformed keys $\phi(\bm{k})$ scaled by $\phi(\bm{k})$, the product of which is multiplied with the transformed queries $\phi(\bm{q})$. The linear map $\phi$ is composed from weight matrices and component functions. The linear map has $s$ components with $a$ being the component weight (see Equation 11). The component weight $a$ can either be fixed (typically at $a=1$) or learnable as proposed by various kernel works (Liu et al., 2022). Due to the fact that the majority of random feature Transformers (e.g., Likhosherstov et al. (2022)) have $a=1$, we follow this setting for consistent comparison between combinations.
@@
3.1 Weight Matrices
In Spectraformer, the weight matrix $\bm{W}$ which is either an approximator (from the families of (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)). or a learner (FastFoodL (Chowdhury et al., 2022)).
Figure 1: Spectraformer framework based on Equation 11
@@ 3.1 Weight Matrices
  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi & Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$. The random matrix $\bm{W}$ can be constructed more efficiently (either reducing time or space complexity) with FastFood (Le et al., 2013). The random matrix $\bm{W}$ can also be enforced to form geometrical couplings such as orthogonality to reduce approximation error with ORF, SORF (Yu et al., 2016). Both these methods produce valid random matrices.
  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $[0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance. The first approach is QMC (Avron et al., 2016), where $\bm{W}:\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\bm{t}_{i}$ being a low discrepancy sequence. $\Phi^{-1}(\bm{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$. MM (Shen et al., 2017; Liu et al., 2022) improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
@@
Liu et al. (2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance. Due to the significant number of weight matrices in random features literature, we decide to cover only the most prominent ones. Appendix A.3 provides the technical details of the approximators discussed here and those that we left out.
Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\bm{W}$. Yang et al. (2015) introduce several of these parametrization schemes, which are then adapted into the attention setting by Chowdhury et al. (2022). The best performing method is via making the constituent matrices of FastFood learnable, which we term FastFoodL. In general, a weight matrix learner is valid as long as it can implicitly model any distribution. In addition the TrigRF, PosRF has also been applied on FastFoodL. Technical details and other learners we have not covered are explored in Appendix A.4. We now explore existing works which have developed from Rahimi & Recht (2007) (see Equation 9, 10) and how they lead to the development of Spectraformer.
3.2 Component functions
@@
While TrigRF(Rahimi & Recht, 2007) is the base component function and guarantees $\mathcal{K}(\bm{\delta})$ to be close to $\phi_{TrigRF}(\bm{x})\cdot\phi_{TrigRF}(\bm{y})$ due to uniform convergence, it is typically evaluated using only the real component of $\exp(-i\bm{\omega}_{j}^{T}\bm{x})$. This leads to $\cos$ and $\sin$ functions which can produce unstable behavior, especially when $K(\bm{\delta})$ is close to 0, where variance is infinite (Choromanski et al., 2021; Likhosherstov et al., 2022). Therefore, TrigRF is avoided in Spectraformer.
  • PosRF (Choromanski et al., 2021) is introduced in Spectraformer as a simple modification over TrigRF to counter this instability. PosRF guarantees positive function output and its variance approaching 0 as $K(\bm{\delta})$ approaches 0. Likhosherstov et al. (2022) generalize TrigRF and PosRF to GERF.
  • OPRF (Likhosherstov et al., 2022) is one of the solution to the reduction of the variance equation of GERF. OPRF is both positive and bounded (whereas PosRF is positive and unbounded). Likhosherstov et al. (2023) generalize GERF to dense matrix instead of scalar parameters leading to DERF.
  • SADERF (Likhosherstov et al., 2023) is one of the solutions to the reduction of the variance equation of DERF. Whilst DERF has many solutions, they all rely on operations which are incompatible with many deep learning libraries on GPU and TPU like SVD and eigen decompositions. SADERF is an extension of GERF with a tighter variance bound on DERF.
@@
The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian. Likhosherstov et al. (2023) and Likhosherstov et al. (2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF. Technical details and other component function(s) not covered are explored in Appendix A.5.

### v5

@@ 2.2. Linearizing Kernelized Attention via Random Features
(7) $\begin{array}[]{c}\bm{A}_{i}=\frac{\sum_{j}\phi(\bm{q}_{i})\phi(\bm{k}_{j})^{T}\bm{v}_{j}}{\sum_{l}\phi(\bm{q}_{i})\phi(\bm{k}_{l})^{T}}=\frac{\phi(\bm{q}_{i})\sum_{j}\phi(\bm{k}_{j})\otimes\bm{v}_{j}}{\phi(\bm{q}_{i})\sum_{l}\phi(\bm{k}_{l})^{T}}\end{array}$
@@
[…]
@@
Oftentimes, however, we might not want to specify a kernel, since kernel choice is indeed a hyperparameter and the softmax is by no means essential or necessary as shown by Tsai et al. (2019). Then kernel learning would be a more suitable option, the detail of which is discussed in Section 2.4. Notwithstanding, we first introduce the main kernel approximation technique: random features.
2.3. Random Features for Kernel Approximation
Of the kernel approximation family of methods, the most successful and appropriate for linearized attention in Transformer has been the random features approach (Rahimi and Recht, 2007). Estimating kernels using random features relies on a fundamental insight from harmonic analysis called Bochner’s theorem (Rudin, 1990) stated as follows:
@@ 2.3. Random Features for Kernel Approximation
(9) $\mathcal{K}(\bm{\delta})=\int p(\bm{\omega})\exp(i\bm{\omega\delta})\mathrm{d}\bm{\omega}$
@@
(10) $\begin{array}[]{ll}\mathcal{K}(\bm{\delta})&=\mathbb{E}[\phi_{\bm{\omega}}(\bm{x})^{T}\phi_{\bm{\omega}}(\bm{y})]\approx<\phi_{\bm{\omega}}(\bm{x}),\phi_{\bm{\omega}}(\bm{y})>\\ \phi_{\bm{\omega}}(\bm{x})\!&=\!\frac{1}{\sqrt{s}}\big[\exp(-\mbox{i}{\bm{\omega}}^{T}_{1}\bm{x}),\ldots,\exp(-\mbox{i}{\bm{\omega}}^{T}_{s}\bm{x})]^{T}\end{array}$
2.4. Random Features for Kernel Learning
Previous random feature techniques are robust in approximating kernels for specific learning problems. However, kernels may be chosen using heuristics and by convention from a popular subset. The kernel is very much a hyperparameter. This could pose a problem since there is no free lunch in learning as much as in kernel choice (Schölkopf and Smola, 2018). Tsai et al. (2019) have shown experimental results that this is also true in the context of attention in Transformer. Hence, instead of picking a kernel for the attention, we could learn the kernel via kernel learning an established kernel methods technique. Kernel learning has shown to be effective in general (Tompkins et al., 2019; Wilson and Adams, 2013; Oliva et al., 2016) and in the case of attention via Gaussian Mixture Model (GMM), learnable weight matrices (FastFood), deep generative model (DGM) (Chowdhury et al., 2022). Kernel learning is achieved by making the weight matrix $\bm{W}$ learnable where $\bm{\omega}_{i}$ is the output of a function parameterizing $p(\bm{\omega})$.
The kernelized attention via random features discussed here is a popular approach, however it is by no means the only one. A more detailed discussion is offered in related work in Section 5. Specifically, Section 5.1 provides alternative kernelized attention formulations and Section 5.2 provides alternative kernel approximations to random features.
Figure 1. Spectraformer framework based on Equation 11
3. Spectraformer
@@
(11) $\begin{array}[]{rl}\phi_{\bm{\omega}}(\bm{x})=&\frac{1}{\sqrt{s}}\big[a_{1}f_{1}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{1}(\bm{\omega}_{s},\bm{x}),...,\\ &a_{1}f_{l}(\bm{\omega}_{1},\bm{x}),...,a_{s}f_{l}(\bm{\omega}_{s},\bm{x})]^{\!\top}\\ \bm{W}=&[\bm{\omega}_{1},...,\bm{\omega}_{s}]^{T}\in\mathbb{R}^{s*d}\end{array}$
Spectraformer is shown in Figure 1. The attention of the OT (on the left-hand side) is replaced by the attention of the Spectraformer which pre-computes the Hadamard (element-wise) product of values $\bm{v}$ and transformed keys $\phi(\bm{k})$ scaled by $\phi(\bm{k})$, the product of which is multiplied with the transformed queries $\phi(\bm{q})$. The linear map $\phi$ is composed from weight matrices and component functions. The linear map has $s$ components with $a$ being the component weight (see Equation 11). The component weight $a$ can either be fixed (typically at $a=1$) or learnable as proposed by various kernel works (Liu et al., 2022). We set the component weight a=1 to ensure a consistent and fair comparison with prior random feature Transformers (e.g., (Likhosherstov et al., 2022)).
@@
3.1. Weight Matrices
In Spectraformer, the weight matrix $\bm{W}$ is either an approximator (from the families of Monte Carlo sampling (e.g., ORF and SORF (Yu et al., 2016)), Unit-cube sampling (e.g., QMC (Avron et al., 2016), MM (Shen et al., 2017)), Quadrature (e.g., SGQ (Dao et al., 2017)). or a learner (e.g., FastFoodL (Chowdhury et al., 2022)).
@@ 3.1. Weight Matrices
  • Monte Carlo sampling-based weight matrix involves approximating the integral of Equation 9 using Monte Carlo sampling with the solution given by Equation 10 (Rahimi and Recht, 2007) with $\bm{W}$ being the ‘Base’ random matrix of $p(.)$, $\bm{\omega}_{i}\sim p(.)$. There are additional methods which improve over Base: structural and geometric methods. Structural methods decompose the Base random matrices $P$ into smaller matrices which reduce time and/ or space complexity. Structural matrices includes FastFoodF (Chowdhury et al., 2022), SCRF (Feng et al., 2015) and P-Model (Choromanski and Sindhwani, 2016). Geometric methods enforce certain geometrical couplings along some dimensions in the construction of random matrices to reduce the approximation error. Geometric methods include: ROM (Choromanski et al., 2017) (ORF (Yu et al., 2016), SORF (Yu et al., 2016)) and SimRF (Reid et al., 2023). In addition to these methods, Monte Carlo has also been adapted as Random Fourier Signature Features (RFSF) (Tóth et al., 2025) for unbiased estimation of the signature kernel, designed for sequential data. RFSF also has more scalable variants, such as RFSF-DP (Diagonally Projected) and RFSF-TRP (Tensor Random Projection), that use additional dimensionality reduction.
  • Unit-cube sampling-based weight matrix transforms Equation 9 to an integral on the unit cube $0,1]^{d}$, then performs the approximation with uniform and independent points, thus reducing variance. The first approach is QMC (Avron et al., 2016), where $\bm{W}:\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$, $\Phi$ being the cumulative distribution function (CDF) associated with $p(.)$ and $\bm{t}_{i}$ being a low discrepancy sequence. $\Phi^{-1}(\bm{t}_{i})\sim p(.)$ whilst having a lower variance than direct sampling from $p(.)$. MM (Shen et al., 2017; Liu et al., 2022) improves over QMC by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$ on $\Phi^{-1}$.
@@
Spectraformer also allows to learn weight matrices by learning $p(.)$ via parameterizing $\bm{W}$. Yang et al. (2015) introduce several of these parametrization schemes, which are then adapted into the attention setting by Chowdhury et al. (2022). This comes from a long line of work in kernel learning. The approach to the dual objectives separates the kernel learning methods into two-stage and one-stage methods (Liu et al., 2022). Two-stage methods (e.g., (Yu et al., 2015; Wilson and Adams, 2013; Bullins et al., 2018)) solves the dual objective separately: $\bm{W}$ is typically first learned via a kernel alignment scheme (Wang et al., 2015), then we solve for $\bm{v}$. One-stage methods (e.g., (Chowdhury et al., 2022; Tompkins et al., 2019)), on the other hand, solves the dual objective in parallel: $p(.)$ corresponding to $\bm{W}$ is parameterized then $\bm{v}$ is solved typically. It is not plausible to apply kernel alignment (Wang et al., 2015) on Transformer architectures. Therefore, only one-stage methods can be considered in the context of weight matrix learner. Two kernel learning methods, GMM and FastFoodL, have been experimented in the non-Transformer setting (Yang et al., 2015), then adapted into the Transformer setting (Chowdhury et al., 2022). DGM (Deep Generative Model) has also been effective in modeling distributions (Kingma and Welling, 2013). It too has been adapted by Chowdhury et al. (2022) for the Transformer setting. Chowdhury et al. (2022) show the superior performance of FastFoodL over other kernel learning-based techniques, hence we have only experimented with FastFoodL among kernel learning techniques.
Liu et al. (2022) show that in various kernel benchmark datasets (among combination with TrigRF), QMC and ORF have consistently high performance. Due to the significant number of weight matrices in random features literature and page limit, we decide to cover only the most prominent ones. Methods which we will not cover in detail includes: SCRF (Feng et al., 2015), P-Model (Choromanski and Sindhwani, 2016), ROM (Choromanski et al., 2017), SimRF (Choromanski et al., 2017), SSF (Lyu, 2017), SSR (Munkhoeva et al., 2018), GMM (Wilson and Adams, 2013; Oliva et al., 2016), DGM (Kingma and Welling, 2013). We still mention them and include them in the figures where appropriate to motivate future work. Figure 2 shows all weight matrices that are implemented and discussed. We now explore these weight matrices in the order specified above.
Figure 2. Spectraformer weight matrices. Dark blue boxes are technique groupings. Arrow denotes which technique belongs to which grouping. Red boxes are techniques which have been explored in the literature but not in this work. Green boxes are techniques which have been explored in the literature and this work. Light blue boxes are techniques not studied empirically previously.
3.1.1. Base
(Rahimi and Recht, 2007) This is the direct sampling from a distribution $\bm{P}_{ij}\sim p$, corresponding to the kernel $\mathcal{K}$ without any additional adjustment, and in approximating the RBF, $P=G$ where $G$ is the normal distribution.
(12) $W_{Base}=\frac{1}{\sigma}\bm{P}$
3.1.2. FastFoodF
(13) $W_{FastFood_{L}}=\frac{1}{\sigma}\bm{SHG\Pi HB}$
[…]
3.1.3. ORF
Orthogonal RF (Yu et al., 2016) decreases the approximation error significantly compared to Base. This is done via replacing $G$ with a properly scaled random orthogonal matrix. However, generating orthogonal matrices become costly quickly as the number of dimensions increases.
(14) $\bm{W}_{ORF}=\frac{1}{\sigma}\bm{SQ}$
3.1.4. SORF
Structured ORF (Yu et al., 2016) decreases time and space complexity of ORF (from $O(d^{2})$ to $O(d\log{d})$ with almost no extra memory cost) by imposing structure on the orthogonal matrices, inspired by structural methods. SORF is unbiased with large $d$. We replace $\bm{S}$ with $\sqrt{d}$ and $\bm{Q}$ with structured matrix $\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
(15) $\bm{W}_{SORF}=\frac{\sqrt{d}}{\sigma}\bm{HD}_{1}\bm{HD}_{2}\bm{HD}_{3}$
3.1.5. QMC
Quasi-Monte Carlo (Avron et al., 2016) evaluates on a low discrepancy sequence (e.g., Halton, Sobol’ Faure, and Niederreiter) of points instead of random points in Monte Carlo. Although the approximation error is only reduced minimally, QMC has been shown to perform better than MC in high dimensions and does not have undesirable clustering effect. To calculate QMC, we first assume that $p(\bm{x})=\prod^{d}_{j=1}p_{j}(\bm{x}_{j})$ factorizes with respect to the dimensions with $p_{j}(.)$ being a univariate density function. Then we define:
(16) $\Phi^{-1}(\bm{t})=\left(\Phi_{1}^{-1}\left(\bm{t}_{1}\right),\ldots,\Phi_{d}^{-1}\left(\bm{t}_{d}\right)\right)\in\mathbb{R}^{d}$
where $\Phi_{j}$ being the cumulative distribution function of $p_{j}$, $\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{s}\in[0,1]^{d}$ being a low discrepancy sequence, and $\bm{\omega}_{i}=\Phi^{-1}(\bm{t}_{i})$. We can thus transform the integral on $\mathbb{R}^{d}$ in Equation 9 to an integral on the unit cube $0,1]^{d}$ as
The weight matrix then can be defined as:
3.1.6. MM
Moment Matching (Shen et al., 2017; Liu et al., 2022) improves over QMC by removing the undesirable clustering effect and having the same approximation error with less features. This is done by replacing $\Phi^{-1}$ with a moment matching scheme $\widetilde{\Phi}^{-1}$:
[…]
3.1.7. SGQ
Sparse Grid Quadrature (Dao et al., 2017) evolves from GQ. GQ (Gaussian Quadrature) (Dao et al., 2017) assumes that the kernel $k$ factorizes with respect to the dimensions and thus can be approximated by a one-dimensional Gaussian quadrature rule. However the total number of points $s$ scale exponentially with the dimensions. However GQ suffers form the curse of dimensionality. This is alleviated using Smolyak rule resulting in SGQ. Assuming the third-degree SGQ using symmetric univariate quadrature points $\left\{-\hat{p}_{1},0,\hat{p}_{1}\right\}$ with weights $(\hat{a}_{1},\hat{a}_{0},\hat{a}_{1})$, then we have:
3.1.8. FastFoodL
This (see Equation 13) is proposed to be learnable by Yang et al. (2015) with two variations: FSARD (scaling matrix $\bm{S}$ previously sampled from chi-squared is made learnable) and FSGBARD (optimizes the marginal likelihood with respect to the diagonal matrices $\bm{G}$ and $\bm{B}$). It is further adapted by Chowdhury et al. (2022) to make $\bm{S}$ or $\bm{S},\bm{G},\bm{B}$ learnable parameters. In our experiments, we maintain this set up.
Figure 3. Spectraformer component functions. Dark blue boxes are technique groupings. Arrow denotes which technique belongs to which grouping. Red boxes are techniques which have been explored in the literature but not in this work. Green boxes are techniques which have been explored in the literature and this work. Light blue boxes are techniques not studied empirically previously.
3.2. Component functions
@@
The alternatives for component functions allow Spectraformer to approximate the RBF with $p(.)$ corresponding to the Gaussian. Likhosherstov et al. (2023, 2022) show that SADERF-ORF is the SOTA combination followed by OPRF-ORF and PosRF-ORF.
Due to page limits, we will not discuss component functions which are not included in our empirical experiments. These techniques are still included in figures and mentioned where necessary to motivate future work in the area. These include DIRF (Likhosherstov et al., 2022) techniques (PoisRF, GeomRF) and HRF (Choromanski et al., 2022).
3.2.1. CDERF
Complex dense exponential random feature is given in Equation 17, where $\bm{\omega},\bm{x}\in\mathbb{R}^{d}$, $p=\mathcal{N}(0,1)^{d}$, $k\in\{1,2\}$ and $\mathbb{S}_{d_{c}}$ being a set of ddd*d complex symmetric matrices. This is the general formula for most of the component functions we are interested in. The parameters and constraints of different CDERF functions are specified in Table 2:
(17) $f_{DE}^{(k)}(\bm{\omega},\bm{x})=Real(D\exp(\bm{\omega}^{T}\bm{A}\bm{\omega}+\bm{\omega}^{T}\bm{B}^{(k)}\bm{x}+\bm{x}^{T}\bm{C}^{(k)}\bm{x}))$
Table 2. An overview of the parameters and constraints for different functions within the Complex Dense Exponential Random Feature (CDERF) family.
$\bm{A}$ $\bm{B}^{(1)}$ $\bm{B}^{(2)}$ $\bm{C}^{(1)}$ $\bm{C}^{(2)}$ $D$ $\bm{x}$ $\bm{y}$
CDERF $\mathbb{S}_{d_{C}}$ $\mathbb{C}^{d*d}$ $\mathbb{C}^{d*d}$ $\mathbb{R}$ $\bm{x}$ $\bm{y}$
DERF (Likhosherstov et al., 2023) $\mathbb{S}_{d}$ $\mathbb{R}^{d*d}$ $\mathbb{R}^{d*d}$ $\mathbb{R}$ $\bm{x}$ $\bm{y}$
TrigRF (Rahimi and Recht, 2007) $0\bm{I}_{d}$ $i\bm{I}_{d}$ $i\bm{I}_{d}$ $0\bm{I}_{d}$ 11 $\bm{x}$ $\bm{y}$
PosRF (Choromanski et al., 2021) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $\bm{I}_{d} 11$ $\bm{x}$ $\bm{y}$
OPRF (Likhosherstov et al., 2022) $0\bm{I}_{d}$ $1\bm{I}_{d}$ $\bm{I}_{d} 11$ $\bm{x}$ $\bm{y}$
3.2.2. TrigRF
Trigonometric RF (Rahimi and Recht, 2007) is the base RFF implementation.
(18) $\begin{array}[]{c}f_{TrigRF}^{(1)}(\bm{\omega},\bm{x})=\sqrt{2}\cos(\bm{\omega}^{T}\bm{x}+\bm{b})\\ f_{TrigRF}^{(2)}(\bm{\omega},\bm{y})=\sqrt{2}\sin(\bm{\omega}^{T}\bm{y}+\bm{b})\\ \bm{b}\sim Uniform(0,2\pi)\end{array}$
3.2.3. PosRF
Positive RF (Choromanski et al., 2021) fixes the problem of TrigRF by enforcing positive component function output in the softmax. The variance of PosRF (in contrast to the variance of TrigRF approaching infinity) approaches 0 as the approximated value of the Softmax kernel approaches 0. PosRF has two forms: Positive RF Base (PosRF-B) (see Equation 19) and Positive RF Hyperbolic (PosRF-Hyp), which is multi-component, i.e., Equation 11: $l=2$, (see Equation 20).
[…]
We only use PosRF-B in our experiments.
3.2.4. GERF
Generalized exponential RF (Likhosherstov et al., 2022) generalizes both TrigRF and PosRF with Equation 21.
[…]
where $\sqrt{.}$ and $\sqrt[n$]{.} denoting a principal root with a complex argument
3.2.5. OPRF
Optimized positive RF (Likhosherstov et al., 2022) is the solution to the minimization of the variance of GERF. Specifically it is defined as Equation 21 with $s=+1$, $\bm{x}+\bm{y}||^{2}>0$, and $\bm{A}\in\mathbb{R}$ defined in terms of $p^{*}$ as:
[…]
3.2.6. DERF
Dense-exponential random features (Likhosherstov et al., 2023) extends GERF and replace $\bm{A},\bm{B},\bm{C}$ with dense matrices. DERF is CDERF when $\bm{B},C$ are in the real instead of the complex plane. Minimizing the variance of DERF leads to two approaches: ADERF and SDERF. However both these approaches rely on SVD and eigen decompositions which are not extensively supported on GPU and deep learning libraries. Therefore SADERF is proposed.
3.2.7. SADERF
Simplified ADERF (Likhosherstov et al., 2023) is a special case of ADERF and extends GERF, requiring only basic unary operations in addition.
(23) $\begin{array}[]{c}f_{SADE}^{(1)}(\bm{\omega},\bm{x})=f_{GE}^{(1)}(\bm{\omega},\Psi\bm{x}),f_{SADE}^{(2)}(\bm{\omega},\bm{y})=f_{GE}^{(2)}(\bm{\omega},\Psi^{-1}\bm{y})\\ \Psi_{l,l}^{*}=(\sum_{j}(\bm{y}_{l}^{(i)})^{2}/\sum_{i}(\bm{x}_{l}^{(i)})^{2})^{1/4}\end{array}$